Pease help me get this answer for this mathe problem? I is on substition Algebra 1 x-y=10
5x-y=-6

Answer by stanbon(48568)

x-y=10
5x-y=-6
Solve the 1st equation for x to get:
x=y+10
Substitute into the 2nd equation to get:
5(y+10)-y=-6
5y+50-y=-6
4y=-56
y=-14
--------
substitute that into the 1st equation to get:
x+14=10
x=-4
---------
Solution: x=-4, y=-14
Cheers,
Stan H.